OpenStudy (anonymous):

Name the y value of the solution to the system of equations. Use substitution: x = 2y - 3 3x + y = 12 (I WILL FAN AND MEDAL)

OpenStudy (anonymous):

Plug the first equation in for x like this 3(2y-3)+y=12 Then solve for x when you get the x valuenplug the x value in for x in the problem to get your y value

OpenStudy (anonymous):

x = 3 & y = 3

OpenStudy (er.mohd.amir):

put x value in second 3 (2y-3)+y=12 6y-9+y=12 7y=12+9 7y=21 y=21/7=3 x=2y-3 x=2*3-3=6-3=3